OpenStudy (anonymous):

How many nanoseconds in 1 year?

OpenStudy (fifciol):

\[1 s=10^{9}ns\] \[1 year=365days=365*24hours=365*24*60 \min=365*24*60*60s\] 365*24*60*60 * 10^9 = 31536 * 10^12 ns

OpenStudy (anonymous):

one year=365*24*60*60=\[3.1536\times10^{7}\]s one second =\[10^{9}\] nanosecond one year=\[3.1536\times10^{7}\times10^{9}\]ns one year=\[3.1536\times10^{16}\]ns

OpenStudy (anonymous):

what you wanna do with that?